OpenStudy (anonymous):

I need an solved exercise for a Statically Indeterminate system.The easiest one you can invent and solve is perfect! :)

OpenStudy (dumbcow):

Any indeterminate system will have more than one solution. The easiest example would be 2 identical equations, leading to an infinite solutions \[2x +3y = 1\] \[4x+6y = 2 \] This system is indeterminate because 2nd equation is just a multiple of 1st. Thus there are infinite solutions. Another example of indeterminate systems is one where there are more variables than equations. \[x +y+z = 3\] \[3x - y +z = 1\] Solutions must be given in terms of an independent variable, leading to multiple solutions. \[x = 1 - \frac{1}{2}z\] \[y = 2 -\frac{1}{2}z\]